# When Is Diversity Rewarded in Cooperative Multi-Agent Learning?

### v1

@@ Abstract
The success of teams in robotics, nature, and society often depends on the division of labor among diverse specialists; however, a principled explanation for when such diversity surpasses a homogeneous team is still missing. Focusing on multi-agent task allocation problems, our goal is to study this question from the perspective of reward design: what kinds of objectives are best suited for heterogeneous teams? We first consider an instantaneous, non-spatial setting where the global reward is built by two generalized aggregation operators: an inner operator that maps the $N$ agents’ effort allocations on individual tasks to a task score, and an outer operator that merges the $M$ task scores into the global team reward. We prove that the curvature of these operators determines whether heterogeneity can increase reward, and that for broad reward families this collapses to a simple convexity test. Next, we ask what incentivizes heterogeneity to emerge when embodied, time-extended agents must learn an effort allocation policy. To study heterogeneity in such settings, we use multi-agent reinforcement learning (MARL) as our computational paradigm, and introduce Heterogeneous Environment Design (HED), a gradient-based algorithm that optimizes the parameter space of underspecified MARL environments to find scenarios where heterogeneity is advantageous. Experiments in matrix games and an embodied Multi-Goal-Capture environment show that, despite the difference in settings, HED rediscovers the reward regimes predicted by our theory to maximize the advantage of heterogeneity, both validating HED and connecting our theoretical insights to reward design in MARL. Together, these results help us understand when behavioral diversity delivers a measurable benefit.
@@ 1 Introduction
Collective systems—from robot fleets to insect colonies—tend to adopt one of two fundamental structures: a uniform shared blueprint or a set of distinct, specialized roles. In multi-agent learning, this dichotomy is reflected in the design choice between behavioral homogeneity (where all agents share one policy) and heterogeneity (where each agent learns its own) bettini2023hetgppo ; bettini2023snd ; rudolph2021desperate . Although diversity unlocks role specialization and asymmetric information use, it also introduces extra coordination cost, representation overhead, and learning complexity chenghao2021celebrating . This trade-off leads us to ask: under what conditions will heterogeneous agents actually outperform the best homogeneous baseline?
A natural setting to study this question in is multi-agent task allocation, where $N$ agents allocate effort across $M$ concurrent tasks. A homogeneous team is forced to adopt a single allocation vector (e.g., every agent spends $0.75$ of its effort on task A and $0.25$ on task B), whereas a heterogeneous team allows agents to specialize differently. This allocation setting underlies multi-agent problems such as coverage and load-balancing Gupta2017 ; Lowe2017 . In this work, we relate it to heterogeneity in cooperative navigation (Sec. 5), Colonel Blotto games, and level-based foraging environments (App. D) Roberson2006 ; noel2022reinforcementcolonelblotto ; papoudakis2021benchmarking_multilevelforaging ; terry2021pettingzoo .
@@
To study heterogeneity in MARL settings not covered by our theoretical analysis, we develop Heterogeneous Environment Design (HED), a gradient-based co-design algorithm that learns environment parameters $\theta$ in under-specified tasks via backpropagation, leveraging differentiable simulation to find parameterizations that either maximize or minimize $\Delta R$. In our experiments $\theta$ parameterizes the reward structure, i.e., the inner- and outer-aggregators, but any differentiable feature of the environment (e.g., obstacle placements) can be optimized. Maximizing the heterogeneity gain through HED allows us to discover previously unknown types of tasks where heterogeneity is essential; minimizing the gain steers an underspecified task toward regimes where homogeneous policies are sufficient, which may be desirable, e.g., when parameter sharing is preferred for learning efficiency (Sec. 4).
We evaluate the validity of our theoretical insights, and HED, in simulation. To this end, we design single-shot and long-horizon reinforcement learning environments whose reward structure instantiates the kinds of aggregation operators we studied. First, in a continuous and a discrete matrix game, we test reward structures based on all nine possible combinations of $\{\min,\mathrm{mean},\max\}$, and find that the heterogeneity gains that result from the agents’ learned policies match our theoretical predictions: concave outer aggregators and convex inner aggregators benefit heterogeneous teams. Next, we test the same aggregators in an embodied, partially observable Multi-Goal-Capture environment, and find that our theory also transfers to this long-horizon MARL setting. Importantly, we also show that reward structures that maximize heterogeneity are meaningful and practically useful. Finally, we find that the empirical heterogeneity gain disappears as the richness of agents’ observations is increased, recovering the finding that rich observations allow agents with identical policy networks to be behaviorally heterogeneous bettini2023hetgppo ; leibo2019malthusian .
We then turn to HED. Across two parameterizable families of aggregators (Softmax and Power-Sum), we show that, despite running on spatial and long-horizon environments, HED rediscovers the reward regimes predicted by our curvature theory to maximize the heterogeneity gain, both validating HED and confirming the connection between our theoretical insights and MARL reward design (Sec. 5).
@@ 1.1 Related Works
Behavioral heterogeneity—where capability-identical agents learn distinct policies—can markedly improve exploration, robustness, and reward bettini2023hetgppo . Yet heterogeneity reduces parameter sharing and thus sample-efficiency, so a core practical question is when its benefits outweigh that cost. Existing MARL methods typically adopt one of two poles: endowing each agent with its own network, or enforcing parameter sharing so that all agents follow a single policy gupta2017cooperative ; rashid2018qmix ; foerster2018counterfactual ; kortvelesy2022qgnn ; sukhbaatar2016learning . A substantial body of work explores the efficiency–diversity trade-off christianos2021scaling ; fu2022revisiting by interpolating between these extremes—e.g. injecting agent IDs into the observation foerster2016learning ; gupta2017cooperative , masking different subsets of shared weights NEURIPS2024_274d0146 , sharing only selected layers chenghao2021celebrating , pruning a shared network into agent-specific sub-graphs kim2023parameter , or producing per-agent parameters with a hypernetwork tessera2024hypermarl . Further, several methods for promoting behavioral diversity in MARL have been proposed, such as: conditioning agents’ policies on a latent representation wang2020roma , decomposing and clustering action spaces wang2021rode , dynamically grouping agents to share parameters yang2022ldsa , applying structural constraints to the agents’ policies bettini2024controlling , or by additional tasks and intrinsic rewards that maximize diversity chenghao2021celebrating ; jaques2019social ; wang2019influence ; jiang2021emergence ; mahajan2019maven ; liu2023contrastive ; liu2024interaction ; li2024learning . While these studies demonstrate how to obtain diversity, they all presume tasks where heterogeneity is already advantageous; what is still missing—and what we supply here—is a principled characterization of which reward structures create that incentive in the first place.
@@
Classic resource–allocation settings, in which a team must divide finite effort among simultaneous objectives, are a central proving ground for cooperative MARL. In robotics, potential-field and market-based learning are the dominant tools for coverage, exploration, and load-balancing tasks Gupta2017 ; Lowe2017 . Game-theoretic analysis and, recently, MARL, play the same role in discrete counterparts such as Colonel-Blotto contests, where players decide how to spread forces over several “battlefields” Roberson2006 ; noel2022reinforcementcolonelblotto . Embodied benchmarks like level-based foraging are heavily studied in MARL, and expose the tension between uniform and specialized effort allocations papoudakis2021benchmarking_multilevelforaging . The survey of BasarOverview highlights how cooperative performance is governed by the shape of the shared reward and the equilibria it induces. Our contribution sharpens this perspective: we prove that the curvature of nested aggregation operators characterizes when heterogeneous allocations dominate homogeneous ones, and introduce algorithmic tools for further exploring settings where diversity is needed.
@@
Co-design is a paradigm where agent policies and their mission or environment are simultaneously optimized gao2024codesign . Our HED algorithm is related to PAIRED dennis2020emergent , a method which automatically designs environments in a curriculum such that an antagonist agent succeeds while the protagonist agent fails. This makes it so that resulting environments are challenging enough without being unsolvable. Similarly, HED designs environments that are advantageous to heterogeneous teams, while disadvantaging homogeneous teams. The key differences are: (1) the environment designer uses direct regret gradient backpropagation via a differentiable simulator instead of RL; this enables higher efficiency by directly leveraging all the environment gradient data available during collection while preventing RL-related issues identified in subsequent works jiang2021replayguided ; parker-holder2021that such as exploration inefficiency and the need for a reward signal; and (2), the protagonist and antagonist are independent multi-agent teams instead of single agents.
2 Formal Setting: Multi-Agent Task Allocation
@@
Consider a set of $N$ agents and $M$ tasks. Each agent $i\in\{1,\ldots,N\}$ allocates effort among the tasks according to the budget constraints: $r_{i1},r_{i2},\ldots,r_{iM}\geq 0\text{ with }\sum_{j=1}^{M}r_{ij}\leq 1$. We can consider both continuous allocations ($r_{ij}$ can be any real number) and discrete allocations ($r_{ij}$ restricted to some finite set of options), with most results in this work focusing on the continuous case. We collect all agents’ allocations into an $N\times M$ matrix: $\mathbf{A}=[r_{ij}]$ [^1] All results in this work can be extended to the case where $r_{i1},r_{i2},\ldots,r_{iM}\geq B_{min}\text{ and }\sum_{j=1}^{M}r_{ij}\leq B_{max}$ for some arbitrary $B_{min},B_{max}\in\mathbb{R}$..
@@
A homogeneous strategy is one where all agents have the same allocation (i.e., devote the same amount of effort to a given task $j$): $r_{ij}=c_{j}\forall\,i,j$. In this case, the allocation matrix $\mathbf{A}$ consists of identical rows. We define $R_{\mathrm{hom}}=\max_{(c_{1},\ldots,c_{M})\in\Delta_{\!\leq}^{M-1}}\;R\Bigl{(}\mathbf{A}\Bigr{)}$ where $\Delta_{\!\leq}^{M-1}=\{(c_{1},\ldots,c_{M})\mid c_{j}\geq 0,\;\sum_{j}c_{j}\leq 1\}$ is the closed unit simplex. A heterogeneous strategy allows each agent $i$ to choose any $(r_{i1},\ldots,r_{iM})\in\Delta_{\!\leq}^{M-1}$ independently. Then $R_{\mathrm{het}}=\max_{\mathbf{A}\in(\Delta_{\!\leq}^{M-1})^{N}}\;R\Bigl{(}\mathbf{A}\Bigr{)}.$ We define the heterogeneity gain as: $\Delta R=R_{\mathrm{het}}-R_{\mathrm{hom}}.$ This quantity measures how much greater the overall reward can be when agents are allowed to specialize differently across tasks, compared to when they must behave identically. Characterizing when $\Delta R>0$ is our main focus in this work.
In MARL, the value $r_{ij}$ represents the ‘local rewards’ agents receive based on task performance, and the aggregate reward $R(\mathbf{A})$ can represent: (i) the payoff of a one-shot effort-allocation game, (ii) the return or sparse terminal reward of an episode, or (iii) the stepwise reward, giving the discounted return $\sum_{t=0}^{T}\gamma^{t}\,R\bigl{(}\mathbf{A}_{t}\bigr{)}$ for a sequence $\bigl{(}\mathbf{A}_{t}\bigr{)}_{t=1,\ldots T}$ of allocations [^2] We can generalize (iii) by considering a sequence $(R_{t})_{t=1\ldots,T}$ of allocation returns that varies per time step.. A positive gain $\Delta R>0$ implies heterogeneous agents can outperform homogeneous ones if the heterogeneous agents can always attain the optimal allocation; if not (as in our Multi-Goal-Capture experiment), this is evidence of an advantage to heterogeneity, but not a formal guarantee.
App. H contains concrete examples of generalized aggregators. In Sec. 5, we study two MARL environments whose reward structure conforms to our setting: a one-shot task allocation game and an embodied Multi-Goal-Capture environment, which can be seen as an extension of various cooperative navigation environments (see terry2021pettingzoo ). Finally, in App. D, we show how the heterogeneity gain of two well-known environments from the literature–Team Colonel Blotto games noel2022reinforcementcolonelblotto and level-based foraging papoudakis2021benchmarking_multilevelforaging –can be studied using the theory we develop in this work.
@@ 3 Analysis
Focusing on continuous allocations, we ask what properties of aggregators guarantee $\Delta R>0$. We draw on the concept of Schur-convexity. Schur-convex functions can be understood as generalizing symmetric, convex aggregators: every convex and symmetric function is Schur-convex, but a Schur-convex function is not necessarily convex roberts1974convex ; peajcariaac1992convex . Proofs for all results are available in App. F.
@@
Let $N,M\geq 2$, and assume that (i) each task-level aggregator $T_{j}$ is strictly Schur-convex and (ii) the outer aggregator $U$ is coordinate-wise strictly increasing. Then either all admissible optimal homogeneous allocations are trivial, or $\Delta R>0$.
@@
[…]
@@
Sum-Form Aggregators. In App. E, we show that the above results reduce to a simple convexity test for sum-form aggregators: a broad class of aggregators that describes most reward structures we consider in this work. This makes testing whether $\Delta R>0$ a simple computation in many cases.
A core topic of this work is reward design: how can we craft team objectives that either advantage or disadvantage behavioral diversity? To do this, it is helpful to first identify an appropriate search space. Our theoretical analysis enables us to narrow down this search space, and focus on aggregators whose curvature can be parametrized. Many family of aggregator functions $\{\,f_{t}(\cdot)\}_{t\in\mathbb{R}}$ can be parametrized by a scalar $t$ which controls whether the aggregation is Schur-convex or Schur-concave, and how strongly it penalizes (or favors) inequalities among the components. For example, the softmax aggregator $\sum_{i=1}^{N}\frac{\exp\bigl{(}t\cdot r_{ij}\bigr{)}}{\sum_{\ell=1}^{N}\exp\bigl{(}t\cdot r_{\ell j}\bigr{)}}$ is parametrized by its temperature, $t$, transitioning from being strictly Schur-concave when $t<0$ to strictly Schur-convex when $t>0$. We can define a space of reward functions by selecting both the task scores and outer aggregator to be softmax functions: let $T_{j}(\mathbf{A})=\sum_{i=1}^{N}\frac{\exp\bigl{(}t\cdot r_{ij}\bigr{)}}{\sum_{\ell=1}^{N}\exp\bigl{(}t\cdot r_{\ell j}\bigr{)}}\;r_{ij},$ and let $U\bigl{(}T_{1}(\mathbf{a}_{1}),\ldots T_{M}(\mathbf{a}_{m}))=\sum_{j=1}^{M}\frac{\exp\bigl{(}\tau\cdot T_{j}(\mathbf{A})\bigr{)}}{\sum_{\ell=1}^{M}\exp\bigl{(}\tau\cdot T_{\ell}(\mathbf{A})\bigr{)}}\;T_{j}(\mathbf{A})$, where $t,\tau\in\mathbb{R}$ parametrize the inner and outer aggregators, respectively. The heterogeneity gain is then dependent on $t$ and $\tau$. As a case study, we derive lower bounds on $\Delta R$ when $N=M$ in Thm. 3.4. We conjecture the lower bound of Thm. 3.4 is actually the exact value of $\Delta R(t,\tau;N)$, but we were unable to prove this. Fig. 2 (righthand side) plots the heterogeneity gains when $N=M=2$.
Theorem 3.4 (Exact softmax heterogeneity gain for $N=M$).
@@
Tab. 3 contains more examples of aggregation operators parameterized by $t$. These families provide a search space for potential reward functions, allowing us to sweep smoothly from $\Delta R=0$ to $\Delta R>0$ reward regimes. As $t\to\pm\infty$, most such aggregators converge to either $\min$ or $\max$, and often reduce to the arithmetic mean for certain parameter choices, motivating us to ask what the heterogeneity gain is when the outer and inner aggregator belong to the set $\{\min,\text{mean},\max\}$. These aggregators are of special interest, since “$\min$” can be seen as a “maximally” Schur-concave function, “$\max$” can be seen as a “maximally” Schur-convex function, and “mean” is both Schur-convex and Schur-concave. Hence, it is worth asking what the heterogeneity gain is when the outer and inner aggregator belong to the set $\{\min,\text{mean},\max\}$. We derive these gains in two cases: continuous allocations where $r_{ij}\in[0,1]$, and discrete effort allocations where $r_{ij}\in\{0,1\}$. The results are summarized in Fig. 2, lefthand side (formal derivation available in App. F).
@@
$T=\min$ $T=\text{mean}$ $T=\max$ Outer $U=\min$ $\Delta R_{\mathrm{F}}$ 0 0 $(M-1)/M$ $\Delta R_{\mathrm{D}}$ 0 $\lfloor N/M\rfloor/N$ $\mathbf{1}_{\{N\geq M\}}$ Outer $U=\text{mean}$ $\Delta R_{\mathrm{F}}$ 0 0 $(M-1)/M$ $\Delta R_{\mathrm{D}}$ 0 0 $(\min\{M,N\}-1)/M$ Outer $U=\max$ $\Delta R_{\mathrm{F}}$ 0 0 0 $\Delta R_{\mathrm{D}}$ 0 0 0

### v2

# 00footnotetext: Supplementary website: https://sites.google.com/view/hetgps

@@ Abstract
The success of teams in robotics, nature, and society often depends on the division of labor among diverse specialists; however, a principled explanation for when such diversity surpasses a homogeneous team is still missing. Focusing on multi-agent task allocation problems, we study this question from the perspective of reward design: what kinds of objectives are best suited for heterogeneous teams? We first consider an instantaneous, non-spatial setting where the global reward is built by two generalized aggregation operators: an inner operator that maps the $N$ agents’ effort allocations on individual tasks to a task score, and an outer operator that merges the $M$ task scores into the global team reward. We prove that the curvature of these operators determines whether heterogeneity can increase reward, and that for broad reward families this collapses to a simple convexity test. Next, we ask what incentivizes heterogeneity to emerge when embodied, time-extended agents must learn an effort allocation policy. To study heterogeneity in such settings, we use multi-agent reinforcement learning (MARL) as our computational paradigm, and introduce Heterogeneity Gain Parameter Search (HetGPS), a gradient-based algorithm that optimizes the parameter space of underspecified MARL environments to find scenarios where heterogeneity is advantageous. Across different environments, we show that HetGPS rediscovers the reward regimes predicted by our theory to maximize the advantage of heterogeneity, both validating HetGPS and connecting our theoretical insights to reward design in MARL. Together, these results help us understand when behavioral diversity delivers a measurable benefit.
@@ 1 Introduction
Collective systems, from robot fleets to insect colonies, tend to adopt one of two structures: a uniform shared blueprint or a set of distinct, specialized roles. In multi-agent learning, this is reflected in the choice between behavioral homogeneity (all agents behave identically) and behavioral heterogeneity (agents specialize) (Bettini et al., 2023; 2025; Rudolph et al., 2021). Such behavioral diversity can be achieved, e.g., via distinct policies (neural heterogeneity) or shared policies conditioning on diverse inputs, such as agent roles (Leibo et al., 2019). Although diversity unlocks role specialization and asymmetric information use, it also introduces extra coordination cost, representation overhead, and learning complexity (Li et al., 2021). This trade-off leads us to ask: under what conditions will heterogeneous agents outperform the best homogeneous baseline?
A natural setting to study this question in is multi-agent task allocation, where $N$ agents allocate effort across $M$ concurrent tasks. Here, we define effort as an abstract quantity representing the agent’s contribution to a given task (e.g., proximity to a goal, or quantity of a task-specific resource the agent gathered), computed in an environment-specific manner. The focus of our work is behavioral, outcome-based heterogeneity, defined through these efforts: a homogeneous team is one where all agents have the same effort allocations (e.g., every agent allocates $0.75$ of its effort to task A and $0.25$ to task B), whereas a heterogeneous team allows agents to achieve specialized allocations. We relate this abstract effort to environmental rewards in many diverse environments, including cooperative navigation, tag, football (Sec. 5), Colonel Blotto games, and level-based foraging (App. E) (Roberson, 2006; Noel, 2022; Papoudakis et al., 2021; Terry et al., 2021). We ask: what effort-based reward functions require heterogeneous behaviors to be maximized?
@@
Algorithmic Search. To study heterogeneity in MARL settings not covered by our theoretical analysis, we develop Heterogeneity Gain Parameter Search (HetGPS), a gradient-based algorithm that optimizes parameters $\theta$ of underspecified, differentiable MARL environments via backpropagation to find configurations that maximize or minimize the empirical $\Delta R$ (we assume differentiability for training efficiency, but consider non-differentiable environments in App. P). While HetGPS can in principle optimize any differentiable environment feature to influence $\Delta R$, we use it here to explore reward structures exclusively, as a means of verifying and extending our theoretical insights. Maximizing the heterogeneity gain allows us to discover reward functions where behavioral diversity is essential. Minimizing the gain leads us to settings where homogeneous policies are sufficient.
Experiments. We validate our theoretical insights, and HetGPS, in simulation, by evaluating in both single-shot and long-horizon reinforcement learning environments whose reward structure instantiates the kinds of aggregation operators studied. First, in a continuous and a discrete matrix game, we test reward structures based on all nine possible combinations of $\{\min,\mathrm{mean},\max\}$, and find that the heterogeneity gains that result from the agents’ learned policies match our theoretical predictions: concave outer aggregators and convex inner aggregators benefit heterogeneous teams. Next, we test the same aggregators in embodied, partially observable environments: Multi-goal-capture, tag, and football. We find that our theory also transfers to such long-horizon MARL settings, and show that reward structures that maximize heterogeneity are meaningful and practically useful. Finally, we find that the empirical heterogeneity gain disappears as the richness of agents’ observations is increased, recovering the finding that rich observations allow agents with identical policy networks to be behaviorally heterogeneous (Bettini et al., 2023; Leibo et al., 2019).
We then turn to HetGPS. Across two parameterizable families of aggregators (Softmax and Power-Sum), we show that, despite running on embodied environments, HetGPS rediscovers the reward regimes predicted by our curvature theory to maximize the heterogeneity gain, validating both HetGPS and the connection between our theoretical insights and MARL reward design (Sec. 5).
@@ 1.1 Related Works
Behavioral Diversity in MARL. Behavioral heterogeneity, where capability-identical agents learn distinct policies, can markedly improve exploration, robustness, and reward (Bettini et al., 2023). Yet heterogeneity reduces parameter sharing and thus sample-efficiency, so a core practical question is when its benefits outweigh that cost. Existing MARL methods typically adopt one of two poles: endowing each agent with its own network, or enforcing parameter sharing so all agents follow a single policy (Gupta et al., 2017a; Rashid et al., 2020; Foerster et al., 2018; Kortvelesy & Prorok, 2022; Sukhbaatar et al., 2016). A large body of work explores the efficiency–diversity trade-off (Christianos et al., 2021; Fu et al., 2022) by interpolating between these extremes: e.g., injecting agent IDs into the observation (Foerster et al., 2016; Gupta et al., 2017a), masking different subsets of shared weights (Li et al., 2024b), sharing only selected layers (Li et al., 2021), pruning a shared network into agent-specific sub-graphs (Kim & Sung, 2023), or producing per-agent parameters with a hypernetwork (Tessera et al., 2024). Further, several methods for promoting behavioral diversity in MARL have been proposed, such as: conditioning agents’ policies on a latent representation (Wang et al., 2020), decomposing and clustering action spaces (Wang et al., 2021), dynamically grouping agents to share parameters (Yang et al., 2022), applying structural constraints to the agents’ policies (Bettini et al., 2024), or by intrinsic rewards that maximize diversity (Li et al., 2021; Jaques et al., 2019; Wang et al., 2019; Jiang & Lu, 2021; Mahajan et al., 2019; Liu et al., 2023; 2024; Li et al., 2024a). While these studies demonstrate how to obtain diversity, they presume tasks where heterogeneity is advantageous. Our work addresses the orthogonal question of when diversity is beneficial, giving a principled characterization of which reward structures create that incentive in the first place.
@@
Task Allocation. Classic resource–allocation settings, in which a team must divide finite effort among simultaneous objectives, are a central proving ground for cooperative MARL. In robotics, potential-field and market-based learning are the dominant tools for coverage, exploration, and load-balancing tasks (Gupta et al., 2017b; Lowe et al., 2017). Game-theoretic analysis and, recently, MARL, play the same role in discrete counterparts such as Colonel-Blotto contests, where players decide how to spread forces over several “battlefields” (Roberson, 2006; Noel, 2022). Embodied benchmarks like level-based foraging are heavily studied in MARL, and expose the tension between uniform and specialized effort allocations (Papoudakis et al., 2021). The survey of (Zhang et al., 2019) highlights how cooperative performance is governed by the shape of the shared reward and the equilibria it induces. Our contribution sharpens this perspective: we prove that the curvature of nested aggregation operators characterizes when heterogeneous allocations dominate homogeneous ones, and introduce algorithmic tools for further exploring settings where diversity is needed.
@@
Environment Co-design. Co-design is a paradigm where agent policies and their mission or environment are simultaneously optimized (Gao et al., 2024; Amir et al., 2025). Our HetGPS algorithm is related to PAIRED (Dennis et al., 2020), a method which automatically designs environments in a curriculum such that an antagonist agent succeeds while the protagonist agent fails. This makes it so that resulting environments are challenging enough without being unsolvable. Similarly, HetGPS designs environments that are advantageous to heterogeneous teams, while disadvantaging homogeneous teams. The key differences are: (1) the environment designer uses direct regret gradient backpropagation via a differentiable simulator instead of RL; this enables higher efficiency by directly leveraging all the environment gradient data available during collection while preventing RL-related issues identified in subsequent works (Jiang et al., 2021; Parker-Holder et al., 2021) such as exploration inefficiency and the need for a reward signal; and (2), the protagonist and antagonist are independent multi-agent teams instead of single agents.
2 Problem Setting
@@
Consider a set of $N$ agents and $M$ tasks. Each agent $i\in\{1,\ldots,N\}$ allocates effort among the tasks according to the budget constraints: $r_{i1},r_{i2},\ldots,r_{iM}\geq 0\text{ with }\sum_{j=1}^{M}r_{ij}\leq 1$, where $r_{ij}$ is defined as the effort agent $i$ puts into task $j$. Here, “effort” $r_{ij}$ is a scalar input to the reward function representing the agent’s contribution to the task, such as resource allocation (App. E) or realized goal proximity (Sec. 5). We can consider both continuous allocations ($r_{ij}$ can be any real number) and discrete allocations ($r_{ij}$ restricted to some finite set of options), with most results in this work focusing on the continuous case. We collect all agents’ allocations into an $N\times M$ matrix: $\mathbf{A}=[r_{ij}]$ [^1] All results in this work can be extended to the case where $r_{i1},r_{i2},\ldots,r_{iM}\geq B_{min}\text{ and }\sum_{j=1}^{M}r_{ij}\leq B_{max}$ for some arbitrary $B_{min},B_{max}\in\mathbb{R}$..
@@
Homogeneous vs. Heterogeneous Strategies. A homogeneous strategy is one where all agents have the same allocation (i.e., devote the same amount of effort to a given task $j$): $r_{ij}=c_{j}\forall\,i,j$. In this case, the allocation matrix $\mathbf{A}$ consists of identical rows. We define $R_{\mathrm{hom}}=\max_{(c_{1},\ldots,c_{M})\in\Delta_{\!\leq}^{M-1}}\;R\Bigl(\mathbf{A}\Bigr)$ where $\Delta_{\!\leq}^{M-1}=\{(c_{1},\ldots,c_{M})\mid c_{j}\geq 0,\;\sum_{j}c_{j}\leq 1\}$ is the closed unit simplex. A heterogeneous strategy allows each agent $i$ to choose any $(r_{i1},\ldots,r_{iM})\in\Delta_{\!\leq}^{M-1}$ independently. Then $R_{\mathrm{het}}=\max_{\mathbf{A}\in(\Delta_{\!\leq}^{M-1})^{N}}\;R\Bigl(\mathbf{A}\Bigr).$ We define the heterogeneity gain as: $\Delta R=R_{\mathrm{het}}-R_{\mathrm{hom}}.$ This quantity measures how much greater the overall reward can be when agents are allowed to specialize differently across tasks, compared to when they must behave identically. Characterizing when $\Delta R>0$ is our main focus in this work.
MARL extension. In MARL, the effort value $r_{ij}$ represents the contribution of agent $i$ to task $j$ as computed by the environment based on agent $i$’s actions. The aggregate reward $R(A)$ can represent: (i) the payoff of a one-shot effort-allocation game, (ii) the return or sparse terminal reward of an episode, or (iii) the stepwise reward, giving the discounted return $\sum_{t=0}^{T}\gamma^{t}\,R\bigl(\mathbf{A}_{t}\bigr)$ for a sequence $\bigl(\mathbf{A}_{t}\bigr)_{t=1,\ldots T}$ of allocations [^2] To extend this further, our theoretical results hold even if the reward function varies over time, $R_{t}(A_{t})$.. $\Delta R>0$ implies that the best heterogeneous policies outperform the best homogeneous ones. In practice, this is evidence of an advantage to heterogeneity and not a formal guarantee, as learning agents may not always converge to optimal policies.
Examples. App. I contains examples of generalized aggregators. Our framework is flexible, and can be applied to many settings, including ones not ordinarily thought of as “task allocation”: in Sec. 5, we apply it to one-shot allocation games, multi-agent navigation, tag, and football. Furthermore, in App. E, we analyze the heterogeneity gain of two well-known environments from the literature: Team Colonel Blotto games (Noel, 2022) and level-based foraging (Papoudakis et al., 2021).
@@ 3 Analysis
Focusing on continuous allocations, we ask what properties of aggregators guarantee $\Delta R>0$. We draw on the concept of Schur-convexity. Schur-convex functions can be understood as generalizing symmetric, convex aggregators: every convex and symmetric function is Schur-convex, but a Schur-convex function is not necessarily convex (Roberts & Varberg, 1974; Peajcariaac & Tong, 1992). Proofs for all results are available in App. G.
@@
Let $N,M\geq 2$, and assume that (i) each task‐level aggregator $T_{j}$ is strictly Schur-convex and (ii) the outer aggregator $U$ is coordinate-wise strictly increasing. Then either all admissible optimal homogeneous allocations are trivial, or $\Delta R>0$.
@@
Let $N,M\geq 2$. If each task‐level aggregator $T_{j}$ is Schur-concave then $\Delta R\;=\;0.$
@@
Sum-Form Aggregators. In App. F, we show that the above results reduce to a simple convexity test for sum-form aggregators: a broad class of aggregators that describes most reward structures we consider in this work. This makes testing whether $\Delta R>0$ a simple computation in many cases.
Parameterizable Families of Aggregators. A core topic of this work is reward design: how can we craft team objectives that either advantage or disadvantage behavioral diversity? To do this, it is helpful to first identify an appropriate search space. Our theoretical analysis enables us to narrow down this search space, and focus on aggregators whose curvature can be parametrized. Many family of aggregator functions $\{\,f_{t}(\cdot)\}_{t\in\mathbb{R}}$ can be parametrized by a scalar $t$ which controls whether the aggregation is Schur‐convex or Schur‐concave, and how strongly it penalizes (or favors) inequalities among the components. For example, the softmax aggregator $\sum_{i=1}^{N}\frac{\exp\bigl(t\cdot r_{ij}\bigr)}{\sum_{\ell=1}^{N}\exp\bigl(t\cdot r_{\ell j}\bigr)}$ is parametrized by its temperature, $t$, transitioning from being strictly Schur-concave when $t<0$ to strictly Schur-convex when $t>0$. We can define a space of reward functions by selecting both the task scores and outer aggregator to be softmax functions: let $T_{j}(\mathbf{A})=\sum_{i=1}^{N}\frac{\exp\bigl(t\cdot r_{ij}\bigr)}{\sum_{\ell=1}^{N}\exp\bigl(t\cdot r_{\ell j}\bigr)}\;r_{ij},$ and let $U\bigl(T_{1}(\mathbf{a}_{1}),\ldots T_{M}(\mathbf{a}_{m}))=\sum_{j=1}^{M}\frac{\exp\bigl(\tau\cdot T_{j}(\mathbf{A})\bigr)}{\sum_{\ell=1}^{M}\exp\bigl(\tau\cdot T_{\ell}(\mathbf{A})\bigr)}\;T_{j}(\mathbf{A})$, where $t,\tau\in\mathbb{R}$ parametrize the inner and outer aggregators, respectively. $\Delta R$ is then dependent on $t$ and $\tau$. Fig. 2 plots $\Delta R$ when $N=M=2$. As a case study, we derive lower bounds on $\Delta R$ when $N=M$ in Thm. 3.4.
Theorem 3.4 (Softmax heterogeneity gain for $N=M$).
@@
Tab. 3 contains more examples of aggregation operators parameterized by $t$. These families provide a search space for potential reward functions, allowing us to sweep smoothly from $\Delta R=0$ to $\Delta R>0$ reward regimes. As $t\to\pm\infty$, most such aggregators converge to either $\min$ or $\max$, and often reduce to the arithmetic mean for certain parameter choices, motivating us to ask what the heterogeneity gain is when the outer and inner aggregator belong to the set $\{\min,\text{mean},\max\}$. These aggregators are of special interest, since $\min$” can be seen as a “maximally” Schur-concave function, $\max$” can be seen as a “maximally” Schur-convex function, and “mean” is both Schur-convex and Schur-concave. Hence, it is worth asking what the heterogeneity gain is when the outer and inner aggregator belong to the set $\{\min,\text{mean},\max\}$. We derive these gains in two cases: continuous allocations where $r_{ij}\in[0,1]$, and discrete effort allocations where $r_{ij}\in\{0,1\}$. The results are summarized in Fig. 2, lefthand side (formal derivation available in App. G).
@@
$T=\min$ $T=\text{mean}$ $T=\max$ Outer $U=\min$ $\Delta R_{\mathrm{F}}$ 0 0 (M1)/M(M-1)/M $\Delta R_{\mathrm{D}}$ 0 $\lfloor N/M\rfloor/N$ $\mathbf{1}_{\{N\geq M\}}$ Outer $U=\text{mean}$ $\Delta R_{\mathrm{F}}$ 0 0 (M1)/M(M-1)/M $\Delta R_{\mathrm{D}}$ 0 0 $(\min\{M,N\}-1)/M$ Outer $U=\max$ $\Delta R_{\mathrm{F}}$ 0 0 0 $\Delta R_{\mathrm{D}}$ 0 0 0